# 3DMambaIPF: A State Space Model for Iterative Point Cloud Filtering via Differentiable Rendering

### v1

@@ 4.4. Differentiable Rendering and Backpropagation
For enhanced filtering of the edges in noisy point clouds, a differentiable rendering method is devised in 3DMambaIPF, explicitly quantifying the level of noise and enabling backpropagation. To render the point clouds, 3D coordinates of the raw point cloud’s are initially converted into the standard coordinate frame via a projective transformation aligned with the camera pose. Subsequently, each discretized point is expressed as scaled Gaussian densities, yielding the occupancy map, and making the backpropagation available. Through the introduction of a differentiable ray tracing operator, these occupancies are transformed into ray termination probabilities. Ultimately, the rendered image is generated by projecting the volume onto the plane.
As depicted in Figure 5, denoised point cloud and adaptive GT are utilized to produce $K$ pairs of images with fixed camera poses using point rendering technique introduced above. These images are then used to calculate a rendering loss, which constitutes a part of the overall loss function. Since the entire rendering process is differentiable, 3DMambaIPF encourages the denoised point cloud to achieve improved geometric appearance, particularly at the boundaries, throughout the backpropagation process.
4.5. Loss Function
@@
Figure 5. Diagram for rendering loss and backpropagation.
Due to the possibility of overlapping regions within each patch during patch partitioning, we introduce a patch stitching method (Zhou et al., 2022). This method weights input points based on the proximity of each point $p_{i}$ to a reference point $p_{r}$ with a Gaussian distribution. Specifically, the weight $w_{i}$ for each $p_{i}$ is calculated as Eq. 6:
(6) $w_{i}=\frac{\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}{\sum_{i}\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}.$
Taking reference from (de Silva Edirimuni et al., 2023), the support radius $r_{s}$ is set to $r/3$, where $r$ is the patch radius. Then, the reconstruction loss at the $t$-th iteration can be defined as Eq. 7:
(7) ${\mathcal{L}}_{Recon}(t)=\sum_{\hat{p_{i}}\in\hat{\mathcal{P}}}w_{i}\min_{{p_{ti}}\in\mathcal{P}_{t}}{{\lVert p_{ti}-\hat{p_{i}}\rVert}^{2}_{2}},$
@@
On the other hand, the rendering loss focuses on visually aligning the noisy point cloud with the adaptive GT using differentiable rendering techniques. As calculated by Eq. 8, it primarily concentrates on optimizing the boundary regions of the point cloud:
(8) ${\mathcal{L}}_{Render}(t)=\frac{1}{K}\sum_{i=1}^{K}\lvert V_{i}(\mathcal{P}_{t})-V_{i}(\hat{\mathcal{P}})\rvert,$
where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, $K$ is the quantity of rendered views, and $V_{i}$ represents a rendered view in a specific camera position. Loss function of one iteration is a weighted combination of the reconstruction loss and the rendering loss, and the ultimate loss is obtained by aggregating loss contributions across iterations, which are summed up as Eq. 9:
(9) ${\mathcal{L}}=\sum_{t=1}^{T}({\mathcal{L}}_{Recon}(t)+\alpha{\mathcal{L}}_{Render}(t)),$
where $T$ is number of iteration times, and weight $\alpha$ is 0.01 in experiments.
5. Experiments
5.1. Datasets and Implementation
Table 1. Results on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
@@ 5.1. Datasets and Implementation
Figure 6. Visualization Comparisons on PU-Net Dataset of 50K points with 2.5% Gaussian deviation. The coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in blue, while those with high P2F values (noisy) are portrayed in red.
@@
Implementation. The 3DMambaIPF network is trained on NVIDIA A100 GPUs with PyTorch 1.13.1 and CUDA 11.7. The network underwent 100 epochs of training on PU-Net dataset, employing the Adam optimizer with a learning rate of 1 × $10^{-4}$. Comparative experiments are conducted with DMRDenoise (Luo and Hu, 2020), Pointfilter (Zhang et al., 2020), PD-Flow (Mao et al., 2022), ScoreDenoise (Luo and Hu, 2021) and IterativePFN (de Silva Edirimuni et al., 2023). All experimental methods (including 3DMambaIPF) are tested on the NVIDIA GeForce RTX 3090 GPU for consistency across evaluations in this paper.
5.2. Evaluation metrics
Figure 7. Visualization Comparisons on large-scale models from the Stanford 3D Scanning Repository with 5% Gaussian deviation. Coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in green, while those with high P2F values (noisy) are portrayed in purple.
Figure 8. Visualization Comparisons on Paris-rue-Madame dataset.
Table 2. Ablation results of loss function on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF (without view loss) 20.02 4.80 30.19 8.13 32.97 10.05 5.94 2.92 7.70 4.14 9.41 5.37
3DMambaIPF (with view loss) 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
Referring to many prior notable works, we utilize Chamfer Distance (CD) and Point-to-Mesh (P2M) Distance as the evaluation metrics for our study. Notably, CD is delineated in Eq. 10, encompassing the combined sum of two components: the average distances between each point in the predicted point cloud and its nearest point in the GT point cloud, and the average distances between each point in the GT point cloud and its nearest point in the predicted point cloud.
(10) ${\mathcal{L}}_{CD}(\hat{\mathcal{P}},\mathcal{P})=\frac{1}{|\hat{\mathcal{P}}|}\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{p}\in\mathcal{P}}\|\hat{p}-{p}\|+\frac{1}{|\mathcal{P}|}\sum_{{p}\in\mathcal{P}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}\|{p}-\hat{p}\|,$
[…]
(11) ${\mathcal{L}}_{P2M}(\hat{\mathcal{P}},\mathcal{M})={\mathcal{L}}_{P2F}(\hat{\mathcal{P}}$ $,\mathcal{M})+{\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M}),$
${\mathcal{L}}_{P2F}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\hat{\mathcal{P}}|}$ $\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{f}\in\mathcal{M}}d(\hat{p},f),$
${\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\mathcal{M}|}$ $\sum_{{f}\in\mathcal{M}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}d(\hat{p},f),$
where $\hat{\mathcal{P}}$ represents the predicted point cloud, $\mathcal{M}$ represents the mesh of GT point cloud, $|\mathcal{P}|$ signifies the number of points in $\mathcal{P}$, $|\mathcal{M}|$ signifies the number of faces in $\mathcal{M}$, and $d(p,f)$ denotes square distance from point $p$ to face $f$. Please note that due to the typically unequal quantities of faces and points, point-wise calculation of P2M distance is unreasonable. In the visualization of denoising results in Section 5.3 and Section 5.4, we utilize the point-wise P2F distance from Eq. 11 to visualize the denoised point cloud results.
Table 3. Ablation results of views in point rendering on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Rendered Views 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF 8 20.36 4.96 30.73 8.48 34.35 10.96 6.04 2.99 8.06 4.36 10.39 6.05
3DMambaIPF 16 20.29 4.97 30.35 8.24 33.08 10.21 6.21 3.14 7.87 4.29 9.63 5.60
3DMambaIPF 24 19.96 4.82 30.26 8.14 32.85 9.95 5.96 2.97 7.81 4.26 9.56 5.53
3DMambaIPF 32 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
Table 4. Ablation results of Mamba layers on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Mamba Layers 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF 1 23.35 6.76 33.24 10.91 36.42 13.17 8.15 4.71 10.82 6.53 13.23 8.20
3DMambaIPF 3 20.24 5.02 30.59 8.61 33.44 10.60 6.12 3.09 8.05 4.43 9.58 5.54
3DMambaIPF 6 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
3DMambaIPF 9 20.55 5.11 30.77 8.59 33.82 10.70 6.15 3.08 8.01 4.38 9.89 5.74
3DMambaIPF 12 20.48 5.05 30.83 8.80 33.92 10.98 6.12 3.07 8.09 4.42 9.85 5.67
Table 5. Ablation results of patch size on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Patch Size 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN 1000 20.55 5.01 30.43 8.43 33.53 10.46 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF (Ours) 1000 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
IterativePFN 2000 20.64 5.02 30.43 8.41 33.93 10.72 6.01 2.99 7.79 4.19 10.25 5.95
3DMambaIPF (Ours) 2000 19.98 4.81 30.05 8.14 32.73 10.04 5.92 2.92 7.52 4.05 8.98 5.13
5.3. Comparisons on Small-Scale Datasets
[…]
5.4. Comparisons on Large-Scale Datasets
The visualization performance of large-scale point clouds is tested on both large-scale synthetic models from the Stanford 3D Scanning Repository and Paris-rue-Madame dataset. In the visualization results of large-scale synthetic datasets, some of the baselines struggle to filter due to the presence of dense and distant noise from the surface. While PointFilter achieves filtering results where each point is close to a clean surface, it lacks significant geometric details and accurate edge structures. It is found that the incorporation of the generative model into PD-Flow results in varied denoising outcomes across different experiments on the same point cloud. Despite potential variations in outcomes, PD-Flow generally falls short in effectively denoising large-scale point clouds, particularly in densely noisy areas where it struggles to achieve significant noise reduction. ScoreDenoise, employing a patch partitioning method, exhibits gaps in results when processing large-scale point clouds. Although IterativePFN addresses this issue with patch-stitching, its performance suffers when handling point cloud edges and densely cluttered point clouds. On the contrary, the inclusion of the Mamba module enhances the processing capabilities of 3DMambaIPF when dealing with dense point clouds. Additionally, the utilization of rendering loss results in remarkable denoising effects specifically on the edges of the point cloud. In the visualization of real-world large-scale dataset (Paris-rue-Madame), when there are no additional noises introduced, all the baseline methods are capable of achieving satisfactory filtering results. 3DMambaIPF outperforms the others in terms of preserving geometric edges.
6. Ablation Study
@@
Ablation experiments are conducted on the influencing factors of 3DMambaIPF, focusing on four aspects: i) loss function, ii) the quantity of rendered images in point rendering, iii) the number of Mamba layers, and iv) patch size. All the ablation experiments are tested on PU-Net dataset with NVIDIA GeForce RTX 3090 GPU. Initially, Table 2 validates the enhancement of the filtering capability of 3DMambaIPF across various levels of noise with the introduction of view loss. Following the previous validation step, we proceed to examine how varying numbers of rendered views affect filtering effectiveness. Table 3 suggests that generally, a greater number of views correlates with stronger denoising ability, with 32 views yielding the best results. However, employing an excessive number of views will decrease the training speed. Subsequently, ablation experiments are carried out on the number of Mamba layers. Table 4 illustrates that the denoising capability is at its peak with 6 Mamba layers. At last, the influence of patch size on denoising capability is examined in Table 5, finding that larger patch sizes are more suitable for larger scenes.
7. Conclusion
@@
In this study, we introduce 3DMambaIPF, an iterative point cloud filtering model that utilizes the Mamba module. 3DMambaIPF not only achieves state-of-the-art results on small-scale (PU-Net) dataset, but also demonstrates superior performance on large-scale datasets, indicating its capability to address denoising challenges in extensive scenes, as validated through ablation experiments. Although 3DMambaIPF may encounter reverse-amplification issues on denoising a smooth and clean surface, its performance remains optimal.

### v2

@@ Introduction
For enhanced filtering of the noisy points around the surface, we propose a differentiable rendering method and integrate it in 3DMambaIPF to explicitly quantify the level of noise and enable backpropagation. To render the point clouds, the 3D coordinates of the raw point clouds are first converted into the standard coordinate frame via a projective transformation aligned with the camera pose. Subsequently, each discretized point is expressed as scaled Gaussian densities, yielding the occupancy map, and making the backpropagation available. Through the introduction of a differentiable ray tracing operator, these occupancies are transformed into ray termination probabilities. Ultimately, the rendered image is generated by projecting the volume onto the plane.
Denoised point cloud and adaptive GT are utilized to produce $K$ pairs of images with fixed camera poses using point rendering technique introduced above. Our differentiable rendering method employs three projection planes, allowing for the rendering of images with greater detail and coverage of the overall shapes, rather than projecting them onto a single plane.
The rendering loss is defined in Equation 1. It aims to characterize the local distribution of points and penalize noise distribution, encouraging the denoised point cloud to achieve an improved geometric appearance, particularly around the surfaces of the point cloud, throughout the backpropagation process:
${\mathcal{L}}_{Render}(t)=\frac{1}{K}\sum_{i=1}^{K}\lvert V_{i}(\mathcal{P}_{t})-V_{i}(\hat{\mathcal{P}})\rvert,$ (1)
where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, $K$ is the quantity of rendered views, and $V_{i}$ represents a rendered view in a specific camera position.
Loss Function
@@
Due to the possibility of overlapping regions within each patch during patch partitioning, we introduce a patch stitching method (Zhou et al. 2022). This method weights input points based on the proximity of each point $p_{i}$ to a reference point $p_{r}$ with a Gaussian distribution. Specifically, the weight $w_{i}$ for each $p_{i}$ is calculated as:
$w_{i}=\frac{\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}{\sum_{i}\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})},$ (2)
where the support radius $r_{s}$ is set to $r/3$, and $r$ denotes to the patch radius. Then, the reconstruction loss at the $t$-th iteration can be defined as:
${\mathcal{L}}_{Recon}(t)=\sum_{\hat{p_{i}}\in\hat{\mathcal{P}}}w_{i}\min_{{p_{ti}}\in\mathcal{P}_{t}}{{\lVert p_{ti}-\hat{p_{i}}\rVert}^{2}_{2}},$ (3)
@@
Loss function of one iteration is a weighted combination of the reconstruction loss and the rendering loss, and the ultimate loss is obtained by aggregating loss contributions across iterations, which are summed up as:
${\mathcal{L}}=\sum_{t=1}^{T}({\mathcal{L}}_{Recon}(t)+\alpha{\mathcal{L}}_{Render}(t)),$ (4)
where $T$ is number of iteration times, and weight $\alpha$ is 0.01 in experiments. The detailed pseudocode for training 3DMambaIPF is provided in Appendix.
Experiments
Figure 5: Visualization Comparisons on Stanford 3D Scanning Repository with 5% Gaussian deviation. Coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in green, while those with high P2F values (noisy) are portrayed in purple.
Figure 6: Visualization Comparisons on PU-Net Dataset of 50K points with 2.5% Gaussian deviation. The coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in blue, while those with high P2F values (noisy) are portrayed in red.
Method Armadillo Dragon
173K points 438K points
CD P2M CD P2M
Noisy 50.19 47.96 38.92 37.51
Pointfilter 6.41 4.82 12.10 12.70
PD-Flow 52.30 50.08 51.48 50.00
ScoreDenoise 29.42 27.41 36.25 34.88
IterativePFN 17.04 14.93 20.43 18.98
3DMambaIPF (Ours) 17.14 15.06 19.79 18.36
Table 1: Results on Stanford 3D Scanning Repository. CD and P2M distances are multiplied by $10^{5}$.
Overview of Comparative Baselines
The filtering capabilities of comparative baselines are conducted with DMRDenoise (Luo and Hu 2020), Pointfilter (Zhang et al. 2020), PD-Flow (Mao et al. 2022), ScoreDenoise (Luo and Hu 2021) and IterativePFN (de Silva Edirimuni et al. 2023). IterativePFN and 3DMambaIPF work on highly-dense and large-scale datasets. Pointfilter tends to obtain over-smoothed surfaces that deviate from the GT and fail to keep the original structure after filtering and struggles to work on highly-dense point clouds. On the other hand, the point cloud filtering task is not addressed as a multi-modal issue in previous works. All the baselines focus solely on either the movement of individual noise points or the generation of clean points, which leads to cluttered visual results, particularly around the surface of the point clouds.
Datasets and Experimental Settings
Details about the PU-Net (Yu et al. 2018) dataset can be found in Appendix. Due to the limited number of points in PU-Net dataset, which contains a maximum of 50K points, point clouds from the Stanford 3D Scanning Repository (Curless and Levoy 1996; Krishnamurthy and Levoy 1996) are utilized to evaluate the filtering performance of large-scale synthetic objects. The preprocessing of datasets is detailed in the Appendix.
The experiments are conducted on NVIDIA A100 GPUs, and 3DMambaIPF is implemented with PyTorch 1.13.1 and CUDA 11.7. The Mamba modules are applied with 6 layers, and the total parameters of Mamba-Denoising Modules is 0.81M. Following previous works (Luo and Hu 2021; de Silva Edirimuni et al. 2023), the quantitative indicators CD and P2M are adopted for evaluation and comparison. More details can be found in Appendix.
@@
Table 2: Results on PU-Net Dataset. CD and P2M distances are multiplied by $10^{5}$.
Rendering Views Mamba Layers 50K points
@@
CD P2M CD P2M CD P2M
/ / 6.05 3.02 8.03 4.36 10.15 5.88
/ 6 5.94 2.92 7.70 4.14 9.41 5.37
8 6 6.04 2.99 8.06 4.36 10.39 6.05
16 6 6.21 3.14 7.87 4.29 9.63 5.60
24 6 5.96 2.97 7.81 4.26 9.56 5.53
32 6 5.89 2.91 7.55 4.05 9.28 5.31
32 1 8.15 4.71 10.82 6.53 13.23 8.20
32 3 6.12 3.09 8.05 4.43 9.58 5.54
32 6 5.89 2.91 7.55 4.05 9.28 5.31
32 9 6.15 3.08 8.01 4.38 9.89 5.74
32 12 6.12 3.07 8.09 4.42 9.85 5.67
Table 3: Ablation results of Mamba layers and rendered views on PU-Net 50K dataset. CD and P2M distances are multiplied by $10^{5}$.
Comparisons on Highly-dense and Large-scale Datasets
The filtering performance of highly dense and large-scale point clouds is evaluated on synthetic models from the Stanford 3D Scanning Repository. The quantitative and visual results are shown in Table 1 and Figure 5. Due to the presence of dense and distant noise relative to the surface, some baselines struggle to effectively filter the point clouds. While Pointfilter achieves the best quantitative results, it lacks significant geometric details and accurate surface structures. It is found that the incorporation of the generative model into PD-Flow results in varied denoising outcomes across different experiments on the same point cloud. Despite potential variations in outcomes, PD-Flow generally falls short in effectively denoising highly-dense point clouds, particularly in severely noisy areas where it struggles to achieve significant noise reduction. ScoreDenoise, employing a patch partitioning method, exhibits gaps in results when processing highly-dense and large-scale point clouds. Although IterativePFN tackles this issue with patch-stitching and achieves competitive quantitative results, its performance is limited when dealing with point cloud surfaces and densely cluttered areas. As a single-modal filtering method, it struggles with the visual quality of point cloud surfaces, leading to gaps on the object surface that do not align with GT. On the contrary, 3DMambaIPF achieves the best performance when dealing with highly dense and large-scale point clouds. Additionally, the utilization of rendering loss results in significantly improved filtering effects and better preservation of geometric details, particularly around the surfaces of the point clouds.
Comparisons on PU-Net Dataset
@@
Ablation Study
Ablation on Modules
@@
Ablation experiments are firstly conducted on the influencing factors of 3DMambaIPF, focusing on two aspects: i) the quantity of rendered images in point rendering, ii) the number of Mamba layers. Table 3 validates the enhancement of the filtering capability of 3DMambaIPF across various levels of noise with the introduction of view loss, and proceeds to examine how varying numbers of rendered views affect filtering effectiveness. However, employing an excessive number of views will decrease the training speed, and detailed runtime metrics will be validated in appendix. Subsequently, ablation experiments are carried out on the number of Mamba layers. Table 3 illustrates that the denoising capability is at its peak with 6 Mamba layers.
Runtime of Differentiable Rendering
An additional ablation for the runtime of rendered views is conducted as shown in Table 4. RuntimeDR represents the total time required to render both the GT and the predicted point clouds. RuntimeStep denotes the total time for one training step. Due to our devised network-free differentiable rendering, the runtime of differentiable rendering only utilizes 11.8%-16.5% of the total time for one training step. Except for this efficiency, differentiable rendering demonstrates notable improvements of 12.0% and 13.9% in CD and P2M metrics, respectively, as the number of views increases. This trade-off between runtime and performance is cost-effective.
Views Metrics
CD P2M RuntimeDR (s) RuntimeStep (s)
8 10.39 6.05 0.08 0.74
16 9.63 5.60 0.11 0.82
24 9.56 5.53 0.14 0.97
32 9.28 5.31 0.17 1.03
Table 4: Runtime analysis on differentiable rendering. CD and P2M metrics are multiplied by $10^{5}$, and tested on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
Comparisons 3D Mamba with Other Backbones
To validate the effectiveness of our designed 3D Mamba, ablation studies (Table 5) comparing the Mamba modules with Transformers and MLPs are conducted. For fairness, the number of modules, layers and embedding sizes are consistent with 3DMambaIPF. For intuitive results, differentiable rendering loss is not implemented in this ablation. 3DMambaIPF fulfills the best denoising performance while ensuring fast training speed as well as the smallest model size.
Methods Metrics
CD P2M GFLOPs FPS MParams
with Transformers 14.02 8.56 178.76 2.31 158.00
with MLPs 30.17 21.71 20.78 4.39 4.56
with Mambas (Ours) 9.42 5.37 16.92 3.62 0.81
Table 5: Comparsions on Mamba with other backbones. CD and P2M metrics are multiplied by $10^{5}$, and tested on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
Conclusion
@@
In this study, we introduce 3DMambaIPF, a novel iterative Mamba-based point cloud filtering model that utilizes a differentiable rendering technique for the first time in this area. 3DMambaIPF not only achieves state-of-the-art results on a commonly-used dataset but also demonstrates superior performance on a highly dense and large-scale dataset, indicating its capability to address denoising challenges in extensive scenes, as validated through ablation experiments. Although 3DMambaIPF may encounter reverse-amplification issues on denoising a smooth and clean surface, its performance remains optimal. In future work, we will explore different datasets to address the challenges of filtering real-world point clouds from large scenes.
Acknowledgments
This project was partially supported by the National Natural Science Foundation of China (U2033209); Hong Kong Innovation and Technology Fund (PiH/213/24GS); the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and an OPPO gift fund.